FEL Upgrade Project Weekly Brief
October 9-13, 2000


The staff of the FEL project, the Accelerator Division, and the entire Jefferson Lab community mourn the untimely loss of our colleague, Bill Kozma.  Bill passed away Sunday night at Duke University hospital.  Bill served the laboratory in many capacities as an administrator in the Accelerator Division and as a friend to all of us.


We continued our preparations for the Oct. 23 semi-annual review for the FEL Upgrade Project.  The financial reports have been received for September so that we will be able to present cost-performance data on
the first four months of the project.

On Oct.12-13, the management of the FEL project participated in a workshop on "Scientific Applications of Accelerator-driven Light Sources" in Washington, DC, sponsored by DOE, SURA, BSA, and LBNL.

WBS 3 (Beam Physics)

Integration of UV and IR systems progressed with solidification of the UV bypass as a single-dipole parallel translation with dispersion suppression via dispersion modulation. Geometry and electron beam optics
based on the Revision 1.0 IR driver design were developed and are being fine tuned.  A layout will be generated.

Data derived from a recent emittance algorithm test was (crudely) reduced and suggests electron beam emittances in the backleg of ~10-12 mm-mrad horizontal and ~3 mm mrad vertical. Planning proceeds to refine the measurement algorithm in support of CSR studies.

WBS 4 (Injector)

Completed a 7 hour run at 20MV/m on a 6mm gap with an average current of 90pA. Have run the FET system with a 4mm gap up to 27 MV/m with an average current of 155pA

The film thickness gage for the aperture cesiator will be sealed with VacSeal.  The cesium load is ready for bake.


We prepared a response to KEK for their offer to electropolish all the cavities for the FEL Upgrade cryomodule.  KEK has recently obtained gradients exceeding 30MV/m with a 9 cell 1.3GHz cavity

WBS 6 (RF)

We continued on the installation of components in Zone 3 and 4.  We expect Zone 4 to be ready for low level and high level debugging tests to begin next week.

We have written the requisitions for the HV slave units for injector rf and expect to have that signed off next week.

Work is progressing according to schedule on the klystron and circulator fabrication.  We sent out technical information packages to four potential vendors for the High Voltage Power Supply upgrades.

WBS 8  (I & C)

The bulk of the week was spent on three items: 1. Rework of the User lab #1 laser safety system, 2. Installation of a new 32 channel pico-motor multiplexer, and 3. Installation of system associated with optical BPMs.  In addition to replacing the LSS for lab #1, there were minor changes done to Labs 3, 4, & 6. The high level and low level certification were completed mid week for 4 & 6 and 3 is in progress today. The 32 channel mux is installed but is lacking one cable, this should be here early next week. Six CAT 5 cables (red for pico-motors) have been pulled and
terminated in each lab. This will allow for a maximum 12 motors per lab.  Initially a combination of cables will total the 32 channels available.  The #@ channel sample and hold VME board is here and being assembled.  This will be used for the initial quad pyro detector - Optical Beam Position Monitors. The optics folks are preparing the transport hardware and lens design.  The new preamp boards for the pyros have arrived today; I will assemble them at 3:00 AM this morning.

The Request for Proposal (RFP) for the 3" shielded beam viewers - all 37 pages- is complete and was signed off 10/13/2000. This will be sent out to shops for bid next week. The req for the 3" BPMs is nearly complete.

The mechanical packages for vacuum control hardware are nearly complete.

WBS 9 (Transport)

o We prepared the existing and now obsolete GX drawing package for use by an outside firm as a template to generate drawing packages of the future GQ and the GX.
o We started generating the GW layout, concentrating on the differences with the GX.  Those areas are the saddle coil leads and the field clamps.
o We started making the magnetic model of the GX where we will try to modeling two candidate materials for the Purcell gap pole faces to achieve the uniformity in a window frame magnet.  Quadrupole - QX (3.125" Quad)
o Prototype Fabrication: Aluminum sample of quadrant is delivered and in measurement.
o Budgetary Quote: Everson has a price based on CNC machining of cores but they felt it was too high.  They are pursuing a price using Wire EDM.
o Our radiation expert is calculating the reflected radiation from severely tilted dump faces as a means of spreading the heat load to larger areas so that the technology of our existing 50 kW dump can by used to absorb the Upgrade's 100 kW.
o Regarding the award of the engineering services contract, best and finals were received from the Offerors, and the technical analysis of the responses to questions was integrated into the overall evaluation.
o We took a count of the vacuum hardware for the machine so that the bulk purchases could start.
o We are about to load test the new cartridge design before placing the final order

WBS 10 (Wiggler)

Bids were received for the dispersion section and the award evaluation is in process.

WBS 11 (Optics)

This week, our upgrade efforts concentrated on the following areas:

Prints for the mirror test stand vacuum enclosure are at the check stage,  and after they go through final review, the drawing package will be released for construction.  The O-BPM's mechanical assemblies were received, and cleaned for vacuum use.  One will be installed in the Optics Control Room (OCR) over the weekend.  We also received
custom-built large area pyroelecric quadrant detectors.


We initiated the fall user run by recommissioning a number of the labs Laser Safety System.  We decided to redo the LSS box that had been in use in Lab 1 since that was the original prototype and we felt it was time to bring it to the same generation as the other boxes.  By the end of the week we had re-certified the safety systems in Labs 4 and 6 and were well on the way to full operation of the other systems.

We made measurements of the harmonic power produced by the FEL as a function of detuning.  In addition to confirming the physics this information is needed for lab safety calculations.  We installed an X-ray setup suitable for 15 keV measurements and were running those tests Friday as this was being written.

As this was the first week of operations, we were involved in hot checkout and measurements of FEL performance.  We reinstalled and checked out the drive laser mode-locker electronics.  We repositioned several optical diagnostics in the OCR in preparation of installation of a permanent, purgeable enclosure.  We remounted some optical cavity
mirrors in our new-style, Ni-plated moly mounts.  We also received our new 5% outcoupling mirrors for producing high average power at ~ 1.05 microns.

Privacy and Security Notice